QUESTION ELEVEN


Suppose Zambia is a small, open economy that has a competitive market for soya beans with a


domestic supply curve is, P = 2Qs and the domestic market demand curve is P+ Qd = 120. On the


world market, soya beans sells for $30 a unit (50kg bag). (Hint: 10 hectograms=1000grams)


a. How much hectograms of soya beans does Zambia import?


b. How much revenue will be raised by a tariff of $10 per unit of soya beans?


c. What is the deadweight loss associated with the tariff?


d. Suppose that instead of a tariff, Zambia creates a quota limiting the number of imports


to only 15 units. How much grams of soya beans will be consumed in Zambia? [15 marks]

1
Expert's answer
2022-01-02T18:21:38-0500

a. In equilibrium Qd = Qs and Pd = Ps, so:

2Q = 120 - Q,

Q = 40 units,

P = 2×40 = 80.

If P = 30, then:

Qd = 120 - 30 = 90 units,

Qs = 30/2 = 15 units,

Qi = 90 - 15 = 75 hectograms of soya beans Zambia imports.

b. A tariff of $10 per unit of soya beans will raise total revenue of TR = 10×75 = $750.

c. Imports with tariff are:

M = Qd - Qs = (120 - (30 + 10)) - 40/2 = 60,

The deadweight loss associated with the tariff is:

DWL = 0.5×15×10 = $75.

d. If instead of a tariff, Zambia creates a quota limiting the number of imports

to only 15 units, then 15 + 15 = 30 of soya beans will be consumed in Zambia.